3. The number of seats in the first row of a theatre has 14 seats. Suppose that each row after the first had 2 additional seats. The number of seats in each row forms an arithmetic sequence.

a) Find the number of seats in the 6th row. [2 marks]

b) Find the total number of seats in the first 10 rows. [3 marks]

c) Given the total number of seats in the first n rows is 660. Show

that n2 13n  660 . Hence, find the value of n when the total number of seats is 660. [6 marks]


1
Expert's answer
2021-11-16T11:38:55-0500

The number of seats in the first row of a theatre has 14 seats. Suppose that each row after the first had 2 additional seats. The number of seats in each row forms an arithmetic sequence.


Let "a_n" be the number of seats in the "n"-th row. Then "a_n=14+2(n-1)."


a) It follows that the number "a_6" of seats in the "6"-th row is equal to "14+2\cdot 5=24."


b) Let us use the formula "S_n=\frac{2a_1+d(n-1)}{2}n" for the sum of "n" terms of arithmetic sequence. The total number of seats in the first 10 rows is equal to "S_{10}=\frac{2\cdot 14+2(10-1)}{2}10=230."


c) Let the total number of seats in the first "n" rows is "660." Then we get that"660=\frac{2\cdot 14+2(n-1)}{2}n= (14+(n-1))n=(n+13)n=n^2+13n." It follows that "n^2+13n-660=0." The last equation is equivalent to "(n-20)(n+33)=0," and thus has the roots "n_1=20" and "n_2=-33." Since the number of seats must be positive, we conclude that the value of "n" is equal to "20."